I need help calculating the real wage based on 1965 prices. What has happened to the purchasing power of the minimum wage over this period

Year min. wage Cpi of1965 = 100
1965 1 100
1976 2.65 185.12
2019 14 809.52
Answer #1

Answer

real wage =(nominal wage/CPI of the year)*100

real wage in the base year is the same as nominal wage

real wage in 1965=(1/100)*100=$1

real wage in 1976=(2.65/185.12)*100

=1.43150389

=1.43

======

real wage in 2019=(14/809.52)*100

=1.7294199

=$1.73

The minimum wage in real term increased over the period from $1 to $1.73.